OpenStudy (anonymous):

Kelly wrote the following steps to prove the equation: Given: 2(x – 3) = 5x + 3 Prove: x = –3

OpenStudy (anonymous):

Step Mathematical Statement Justification 0 2(x - 3) = 5x + 3 Given 1 2x - 6 = 5x + 3 Distributive Property 2 – 6 = 3x + 3 Subtraction Property of Equality 3 3x + 3 = –6 Transitive Property of Equality 4 3x = –9 Subtraction Property of Equality 5 x = –3 Division Property of Equality

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

do you understand it right ?